OpenStudy (anonymous):

rachel must store a few pieces of her furniture for six months. a storage unit costs $55 for the first month and $47.50 for the remaining months. How much will it cost to store her furniture?

OpenStudy (anonymous):

HI... \[\Huge \color{red}{\mathbb{WELCOME}}~~\color{orange}{\mathbb{TO}}~~\color{purple}{\mathbb{OPENSTUDY}}\color{blue}{!!}\]

OpenStudy (anonymous):

since she needs to store the furniture for 6 months, and you know she pays $55 for the the 1st month and pays $47.50 for the remaining five months, \[T_c= $55+5($47.50)\] \[T_c=??\]

OpenStudy (anonymous):

T(c) is the total cost she pays for storing the furniture for 6 months..

OpenStudy (anonymous):

do you understand what i have written??

OpenStudy (anonymous):

not really ? this are my answers i hav to choose ? a) $237.50 b) $285.00 c) $292.50 d) $330.00

OpenStudy (texaschic101):

55 + 5(47.50) = your answer can you do this problem ?

OpenStudy (texaschic101):

@kausarsalley ....explained it well

OpenStudy (anonymous):

292.5 ?

OpenStudy (texaschic101):

Bingo....you got it :)

OpenStudy (anonymous):

thanks :) that kind of help me :) i still have alot to do :/ but im not really good in math

OpenStudy (texaschic101):

math is not my best subject either....

OpenStudy (texaschic101):

do you have any more questions ?

OpenStudy (anonymous):

yeahh :/ i need to do alot but what if bother you helping me :/ ?

OpenStudy (texaschic101):

I can try to help you, but remember, math is not my best subject. I will do what I can.

OpenStudy (anonymous):

thanks (: but you can really help me {: ?

OpenStudy (texaschic101):

possibly...what you got ?

OpenStudy (anonymous):

what did i get on the questions ?

OpenStudy (texaschic101):

huh ?....what do you mean ?

OpenStudy (anonymous):

wait are you talking about whats my next qustion ?

OpenStudy (texaschic101):

yes......what other questions do you need help with ?

OpenStudy (anonymous):

George bought 7.5 cases of pancake mix for his dinner. a case contains 12 1.3-pound boxes. How many pounds of pankcake mix did George buy?

OpenStudy (texaschic101):

George bought 7.5 cases and each case contains 12 1.3 lb boxes. Each case contains (12)(1.3) = 15.6 lbs He bought 7.5...so (7.5)(15.6) = 117 lbs. ANSWER : George bought 117 lbs.

OpenStudy (anonymous):

thanks that helped alot i wasnt even close to that answer /:

OpenStudy (texaschic101):

With word problems, you just have to break them down. Sometimes it gets confusing because there is sometimes stuff in word problems that you don't even need to use to solve it.

OpenStudy (texaschic101):

got anymore ?

OpenStudy (anonymous):

Sandy is working with a carpenter to frame a house. They are using 8-foot-long boards, but each board must be cut to 94.6 inches long How much is cut off each board? (1foot= 12 inches)

OpenStudy (texaschic101):

1/12 = 8/x (1 ft to 12 in = 8 ft to x in) cross multiply x = (8)(12) x = 96 in Each 8 ft board is 96 inches. Each board has to be cut to 94.6 in so we have to subtract 96 - 94.6 = 1.4 ANSWER : 1.4 inches are cut off of each board.

OpenStudy (anonymous):

i got it now thanks :) im understanding little by little

OpenStudy (texaschic101):

cool.......if you need any more help...just tag me...good luck

OpenStudy (anonymous):

okayy (: are you going to be online ?

OpenStudy (texaschic101):

for a little while

OpenStudy (anonymous):

well can you help me with these last 4questions ?

OpenStudy (texaschic101):

lets see them :)

OpenStudy (anonymous):

A machine uses pieces of tape to seal packages. It cuts the tape into pieces exactly 2.375 inches long. if it uses 4pieces of tape for a package , how many packages can it seal with a roll of tape 1,900 inches long ?

OpenStudy (texaschic101):

It uses 4 pieces of tape for a package and the tape is 2.375 inches long, then it uses (4)(2.375) = 9.5 It uses 9.5 inches per package. How many packages can it tape using 1900 inches ? We have to divide......1900/9.5 = 200 ANSWER : It can tape 200 packages with 1900 inches.

OpenStudy (anonymous):

thanks :) .. but i kind of feel dumb i cant even do these that well :/

OpenStudy (texaschic101):

The next one we will work on together......step by step.....don't feel dumb.....you are learning.

OpenStudy (anonymous):

Jack"s club is selling grapefruit to raise money. for every box they sell , they get $1.35 profit. They have sold 84 boxes already. How many more boxes must they sell to raise 270 dollars ?

OpenStudy (texaschic101):

First lets see how much profit they have already made. They get 1.35 for every box they sell. They have already sold 84 boxes. Do you know how to tell how much profit they have already made ?

OpenStudy (anonymous):

113.4 ?

OpenStudy (texaschic101):

yes.......now do you know how to do proportions ?

OpenStudy (anonymous):

not really

OpenStudy (texaschic101):

let me show you..... 84 boxes made them 113.40 profit. Now you want to find out how many boxes that will make them 270. 84/113.4 = x/270 (x stands for # of boxes.) This is the proportion that we will have to solve to find out how many boxes will make $270. I have a really hard time explaining proportions. Can you kind of understand ?

OpenStudy (anonymous):

kind of ? so do you add 113.4 ? :o

OpenStudy (texaschic101):

no......we do the proportion... 84/113.4 = x/270 cross multiply (113.4)(x) = (84)(270) 113.4x = 22680 can you solve for x ? Do you know how ?

OpenStudy (anonymous):

no ;/ ?

OpenStudy (texaschic101):

The more I look at this question.......I might have done it wrong. They already made 113.40 selling 84 boxes. So all they need is 270 - 113.40 = 156.60 Is the question asking for a total of $270 or is it $270 more then what they already sold ?

OpenStudy (anonymous):

i got it thanks :) i need one more :/ ?

OpenStudy (texaschic101):

did you figure it out ?

OpenStudy (anonymous):

yeahh .. i have one more question ?

OpenStudy (texaschic101):

ok

OpenStudy (texaschic101):

For the question I messed up on....did you get 116 boxes ?

OpenStudy (anonymous):

a stadium has 20,000 reserved seats and 25,00 general admission seats. if reserved seats are $17.50 and general admission seats $10.75 , how much money would be taken in if tickets for all the seats were sold ?

OpenStudy (texaschic101):

20,000 reserved seats at 17.50 = (20,000)(17.50) = the money you will get for reserved seats. Is the general admission seats = 2,500 or 25,000 ? I need to know this so I can continue with the problem

OpenStudy (anonymous):

25,000

OpenStudy (texaschic101):

ok.... general admission seats....there are 25,000 at 10.75 (25,000)(10.75) = the money you will get for general admission seats. can you figure the money you will get from all seats ?

OpenStudy (anonymous):

no ;/ these are my answers ;/ a) $718,750 b) $268,750 c) $350,000 d) $618,750

OpenStudy (texaschic101):

(20,000)(17.50) = 350,000 (reserved seat money) (25,000)(10.75) = 268,750 (general admission seat money) 350,000 + 268,750 = 618,750 (total money) ANSWER : D. $ 618,750

OpenStudy (texaschic101):

That question that I messed up on.....the answer is 116 boxes

OpenStudy (anonymous):

thanks :) well im going to go eat are you still going to be on later

OpenStudy (texaschic101):

I am not sure....possibly....if I am, I can help....enjoy your lunch :)

OpenStudy (anonymous):

okayy ill message you :) ? and talk to you soon

OpenStudy (texaschic101):

ok...bye for now :)